OpenStudy (anonymous):

what is the simplified form of (3 / 4x+3) + (21 / 8x^2-14x-15) ? Help!

OpenStudy (anonymous):

Ok. For this, you have to factor the \(8x^2 - 14x - 15\). Do you know how to do that?

OpenStudy (anonymous):

not when the first and last numbers dont have factors that end up equaling -14

OpenStudy (anonymous):

Find factors of -120 that add up to -14. How about -20 and 6?

OpenStudy (anonymous):

So: \[8x^2 - 14x - 15 \rightarrow 8x^2 - 20x + 6x - 15 \rightarrow (8x^2 - 20x) + (6x - 15) \rightarrow ?\]

OpenStudy (anonymous):

ohhh okay

OpenStudy (anonymous):

Tell me what you get.

OpenStudy (anonymous):

6(2x+5) / (2x-5)(4x+3)

OpenStudy (anonymous):

Yup. Now back to the original problem. \[\frac{3}{4x + 3} + \frac{21}{(2x - 5)(4x + 3)}\rightarrow \frac{3}{4x + 3} \times \frac{2x - 5}{2x - 5}+ \frac{21}{(2x - 5)(4x + 3)}\]Can you try to do the rest?

OpenStudy (anonymous):

Its all on paper

OpenStudy (anonymous):

its alot is there like a way i can send a picture?

OpenStudy (anonymous):

wait so 6(2x+5) / (2x-5)(4x+3) isnt the final answer??

OpenStudy (anonymous):

No. \[\frac{3}{4x + 3} \times \frac{2x - 5}{2x - 5}+ \frac{21}{(2x - 5)(4x + 3)} \rightarrow \frac{3(2x - 5) + 21}{(2x - 5)(4x + 3)}\]Think you can finish what's in the numerator?

OpenStudy (anonymous):

\[\frac{6x - 6}{(2x - 5)(4x + 3)}\]That's what the end result is.

OpenStudy (anonymous):

Do you get it?

OpenStudy (anonymous):

ya but wouldnt it be 6x+1 / (2x-5)(4x+3) ?

OpenStudy (anonymous):

No. 3(2x - 5) = 6x - 15. THen, there's the + 21. 21 - 15 = 6 See?

OpenStudy (anonymous):

yeah that makes sense but the workbook says a.) 6(2x+5) / (2x-5)(4x+3) b.) 6/ 4x+3 c.) 6x+1 / (2x-5)(4x+3) d.) 6x+1 / (2x+5)(4x+3) are the only options

OpenStudy (anonymous):

I don't see how you can get any of those...

OpenStudy (anonymous):

This is the only thing that I can see possible... \[\frac{6(x + 1)}{(2x - 5)(4x + 3)}\]

OpenStudy (anonymous):

yeah thats what i thought too. Oh well

OpenStudy (anonymous):

Sorry I can't be of more help...

OpenStudy (anonymous):

@ganeshie8 Do you see my mistake anywhere?

ganeshie8 (ganeshie8):

No. looks perfect to me. and wolfrom confirms the same :) http://www.wolframalpha.com/input/?i=+%283+%2F+%284x%2B3%29%29+%2B+%2821+%2F%28+8x%5E2-14x-15%29%29

ganeshie8 (ganeshie8):

must be printing mistake in the options...

OpenStudy (anonymous):

thats so weird! at least i know how to do it, thanks guys :)

OpenStudy (anonymous):

Alright. np :) And thanks ganeshie8